An ingestion tablet contains a mass of 0.30 g of magnesium hydroxide [Mg(OH)2] as its only basic ingredient. The balanced chemical equation between magnesium and hydrochloric acid (HCl(aq)), the acid produced in the stomach is as follows:

          Mg(OH)2(aq) + 2HCl(aq) ⟶ MgCl2(aq) + 2H2O(l)

(i) Calculate the volume of 1.00 M HCl neutralised by two of these ingestion tablets?

(ii) What mass of the salt is formed in this neutralisation reaction?

(iii) How many magnesium ions are present in this amount of a salt?


1
Expert's answer
2022-03-27T10:28:11-0400

(i) Mass of Magnesium hydroxide present in ingestion tablet = 0.30 g

Balanced equation


"Mg(OH)_2(aq)+2HCl(aq) \longrightarrow MgCl_2(aq)+2H_2O( l)"


Molarity of 1M HCl = 1.0 M HCl


moles of "Mg(OH)_2= \frac{mass}{molarmass}=\frac{0.3}{58}=0.00517"


"Molarity = \frac{mole}{Volume}\times 1000"


"1 = \frac{0.00517}{2\times Volume}\times 1000"


Volume = 2.585 g


(ii) Moles of salt formed= 0.00517

mass of salt = "95.2\times 0.00517=" 0.49 g


(iii) moles of Magnesium salt = 0.00517


moles of Magnesium ion = 0.00517

Number of magnesium = "6.023\times 10 ^{23} \times 0.00517 = 3.11 \times 10^{21} ions"